16x squared - 25=0
Answer by lenny460(1073)

16x squared - 25 = 0
I will assume that equation is:
16x^2 - 25 = 0
We will use the Difference of Squares Formula to solve the above:
Factoring:
(4x - 5)(4x + 5) = 0

Answer:
x = 5/4, -5/4







Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com